OpenStudy (anonymous):

Organotin compounds play a significant role in diverse industrial applications. They have been used as plastic stabilizers and as pesticides or fungicides. One method used to prepare simple tetraalkylstannanes is the controlled direct reaction of liquid tin(IV) chloride with highly reactive trialkylaluminum compounds, such as liquid triethylaluminum (Al(C2H5)3). 3SnCl4 + 4Al(C2H5)3 --> 3Sn(C2H5)4 + 4AlCl3 In one experiment, 0.230 L of SnCl4 (d = 2.226 g/mL) was treated with 0.485 L of triethylaluminum (Al(C2H5)3); d = 0.835 g/mL). What is the theoretical yield in this experiment

OpenStudy (abmon98):

Density=Mass(g)/Volume(ml) Change L to ml by * by 1000 Mass/230=2.226 Mass=2.226*230 Mass=511.98 grams of SnCl4 Number of Moles=Mass(g)/Molar Mass(g/mol) N of Mole=511.98/(118.710)+(35.5*4) N of Mole=511.98/260.71 N=1.9637 moles of SnCl4 moles Al(C2H5)3 required=1.9637*4/3=2.618 moles Density=Mass(g)/Volume(ml) Change L to ml by * by 1000 0.835=Mass(g)/485 Mass=485*0.835 Mass=404.975 grams of (Al(C2H5)3) Number of Moles=Mass(g)/Molar Mass(g/mol) 404.975/(26.98+(12*6)+(1*15))=404.975/113.98=3.553 moles of (Al(C2H5)3) moles SnCl4 required=3.553*3/4=2.66 moles We realise that SnCl4 is the limiting Reagent and Al(C2H5)3) is in excess.

OpenStudy (abmon98):

Look at the mole Ratio of SnCl4 to Sn(C2H5)4 they both share the same number of moles 1.9637*234.9544=461.3799

OpenStudy (aaronq):

^no units, \(-1\).. jk haha

OpenStudy (abmon98):

lol :D